OpenStudy (curry):

Assume that 2-3i is a solution of ax^2+bx+c=0, where a,b,c are real numbers, which of the following is also a solution of the equation? show WORK!!! 2+3i -2-3i -2+3i 3+2i 1/(2-3i) Assume that 2-3i is a solution of ax^2+bx+c=0, where a,b,c are real numbers, which of the following is also a solution of the equation? show WORK!!! 2+3i -2-3i -2+3i 3+2i 1/(2-3i) @Mathematics

OpenStudy (anonymous):

@curry this is not any work to show. if \[a+bi\] is a zero of a polynomial with real coefficients then so is \[a-bi\]

OpenStudy (anonymous):

^

OpenStudy (anonymous):

i guess we could say we know this by the quadratic formula. it says the solutions are \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] and so if \[b^2-4ac<0\] you have two complex solutions, \[-\frac{b}{2a}+\frac{\sqrt{4ac-b^2}}{2a}i\] and \[-\frac{b}{2a}-\frac{\sqrt{4ac-b^2}}{2a}i\]

OpenStudy (anonymous):

you can actually show it though using the formula but it's implied

OpenStudy (anonymous):

i wonder if you could solve it by = the one to your answer and then finding a b and c lol

OpenStudy (anonymous):

by using like another formula